OpenStudy (anonymous):

x - y = 9 x + 2y = -6 Choose one answer. a. (5,4) b. (4,-5) c. (4,5) d. (-5,-4)

OpenStudy (anonymous):

x - y = 9 x + 2y = -6 subtract the two equations simulataneously -3y = 15 y = -5 now solve for x

OpenStudy (apoorvk):

solve these two equations! subtract the whole of the first equation from the second! the 'x' gets eliminated, and you can find out 'y' then. And use this value of 'y' to get 'x'!

OpenStudy (anonymous):

Subtract the two equations what you got??

OpenStudy (anonymous):

ohk hold on

OpenStudy (anonymous):

Sure, take your time..

OpenStudy (anonymous):

-3y=-15?

OpenStudy (anonymous):

-3y=15

OpenStudy (anonymous):

-3y = 15 is right. now solve for y

OpenStudy (anonymous):

you do that by dividing both sides by the coefficient of y

OpenStudy (anonymous):

Hi abb50.. She is a good girl.. She will calculate on her own.. Do not need to tell her...

OpenStudy (anonymous):

i didn't tell her anything

OpenStudy (anonymous):

Dividing both sides by......................etc etc..

OpenStudy (anonymous):

it was just a hint. is that not allowed?

OpenStudy (anonymous):

Let her do on her own.. if not then you surely can give hints...

OpenStudy (anonymous):

okay sorry

OpenStudy (anonymous):

lol what? abb50 you have nothing to apologize for.

OpenStudy (anonymous):

I do not want to hear sorry from you abb50.. Even I am sorry if I talked rude to you...

OpenStudy (anonymous):

thnx both of u guys i got the y but i got 14 for the x which im sure is wrong

OpenStudy (anonymous):

@abb50 u were helping nd thnx i got -5 anyway even without u giving it to me lol i didnt even realized u gave me the answer thats how dumb i am!

OpenStudy (anonymous):

nd thanx @waterineyes for all ur help!!! i am learning a lot from u!!!

OpenStudy (anonymous):

No need to say thanks... See, from first equation: x - y = 9 So here, x = 9 + y put y: x = 9 + (-5) Now solve for x...

OpenStudy (anonymous):

ohk hold on

OpenStudy (anonymous):

4!

OpenStudy (anonymous):

so the answer is b?

OpenStudy (anonymous):

That's the one with x=4 and y=-5 ;)

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

thnx all of u!!